OpenStudy (ayee_ciera):

Solve the system by substitution. 2x-y+z=-4 z=5 -2x+3y-z=-10 a.) (-8, 7, 5) b.) (–8, –7, 5) c.) (8, –7, 5) d.)(–8, –7, –5)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

2x-y+z=-4 2(-8)-(-7)+5 -16+7=-9+5=-4